OpenStudy (anonymous):

in each part, use the given inner product on R^2 to find || w|| where w=(-1,3) a)the Euclidean inner product b) the weighted Euclidean inner product = 3u1v1+2u2v2 where u =(u1,v1) and v=(v1,v2) c) the inner product generated by the matrix (top: 1,2; bottom: -1,3) I don't understand the question. Please tell me what they want me to do?

OpenStudy (anonymous):

\[ ||w||=<w,w>=3(w_1)(w_1)+2(w_2)(w_2) \]

OpenStudy (anonymous):

thats part (b)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

par a

OpenStudy (anonymous):

= sqtr(10)

OpenStudy (anonymous):

I'm ok withc

OpenStudy (anonymous):

(c): obtain w' \[w'=\left[\begin{matrix}1&2\\-1&3\end{matrix}\right]\left[\begin{matrix}-1\\3\end{matrix}\right]\]

OpenStudy (anonymous):

a) is it right? I got c) (know how to get)

OpenStudy (anonymous):

let me rearrange mine: a) = sqr(10) . is it right? everybody?

OpenStudy (anonymous):

b) and c) is quite easy, I got it. just want to confirm a

OpenStudy (anonymous):

if you take square-root, it is the L2 norm!

OpenStudy (anonymous):

what if in part b, it's not +, but -? I mean what if the problem ask for finding d(u,v) , base on the inner product they give out <u, v> = 3u1v1+2u2v2